# in Autonomous Driving

### v1

@@ Abstract
Understanding the traffic scenes and then generating high-definition (HD) maps present significant challenges in autonomous driving. In this paper, we defined a novel Traffic Topology Scene Graph ($\text{T}^{2}\text{SG}$), a unified scene graph explicitly modeling the lane, controlled and guided by different road signals (e.g., right turn), and topology relationships among them, which is always ignored by previous high-definition (HD) mapping methods. For the generation of $\text{T}^{2}\text{SG}$, we propose TopoFormer, a novel one-stage Topology Scene Graph TransFormer with two newly-designed layers. Specifically, TopoFormer incorporates a Lane Aggregation Layer (LAL) that leverages the geometric distance among the centerline of lanes to guide the aggregation of global information. Furthermore, we proposed a Counterfactual Intervention Layer (CIL) to model the reasonable road structure (e.g., intersection, straight) among lanes under counterfactual intervention. Then the generated $\text{T}^{2}\text{SG}$ can provide a more accurate and explainable description of the topological structure in traffic scenes. Experimental results demonstrate that TopoFormer outperforms existing methods on the $\text{T}^{2}\text{SG}$ generation task, and the generated $\text{T}^{2}\text{SG}$ significantly enhances traffic topology reasoning in downstream tasks, achieving a state-of-the-art performance of 46.3 OLS on the OpenLane-V2 benchmark. We will release our source code and model.
@@ 1 Introduction
Understanding the traffic scene is the key component of autonomous driving. Except for detecting and recognizing individual elements, the vehicles need to infer the topology relationship among them. Conventional traffic scene understanding tasks, such as lane perception [24, 41, 44], road signal elements detection [6], and high definition (HD) mapping [27, 31] focus primarily on the isolated elements (i.e., map elements, and road signal elements ) but miss the relationship among them. To model the unified HD mapping for lane perception and road signal elements association, the traffic topology reasoning task on OpenLane-V2 [49] has recently been proposed. As a map-like reasoning results shown in Figure 1(c), the reasoning task aims to create a topology graph of the detected elements and thus facilitate decision-making in the downstream tasks, such as ego planning [8, 7] and motion prediction [1, 39].
@@
(2) We propose a Topology Scene Graph Transformer (TopoFormer) for $\text{T}^{2}\text{SG}$ generation task, which captures global dependencies among lanes with a Lane Aggeration Layer. Additionally, we introduce a Counterfactual Intervention Layer to emphasize the reasonable road structure influencing lane connectivity, demonstrating effectiveness in topology reasoning tasks.
(3) We evaluate our TopoFormer in scene graph generation task and outperform all state-of-the-art methods. Furthermore, we attain a 46.3 OLS on the traffic topology reasoning benchmark, OpenLane-V2 [49], demonstrating the effectiveness of the proposed $\text{T}^{2}\text{SG}$ and TopoFormer for downstream tasks.
@@ 2 Related Work
Scene Graph Generation. Scene graphs were first introduced [19] for Image Retrieval, which broke down an image into its constituent objects, their attributes, and the relationships between them. Following this, the advent of the Visual Genome dataset [22] marked a significant advancement in the field of image understanding. This dataset not only enhanced the comprehension of images but also facilitated the evolution of various methodologies for scene graph extraction [58, 11, 18]. An increasing number of studies are exploring the application of scene graphs in downstream tasks, such as image captioning [15, 57] and visual question answering [50, 10].  [48] introduced a 3D scene graph to enhance the perception of relationships between objects in 3D space. Subsequent research explored 3D scene graph generation methods based on GCN and Transformer  [60, 51, 33, 21, 35], and applied in point cloud registration  [43, 54] and scene generation  [14, 59]. Unlike previous methods, $\text{T}^{2}\text{SG}$ is the first to apply scene graphs in traffic scene understanding, which treats lanes that are controlled and guided by different road signals as nodes and considers the connections among lanes as edges.
@@
Figure 2: The overview of our proposed TopoFormer. Given the input multi-view image, we employ a DETR-like detector to identify lane objects with corresponding class and centerline coordinates. Subsequently, TopoFormer infers the relationships among these objects, which, along with the objects themselves, constitute the $\text{T}^{2}\text{SG}$. The main components of TopoFormer include two newly designed layers: (a) the Counterfactual Intervention Layer incorporating Counterfactual Self-Attention, and (b) the Lane Aggregation Layer incorporating Geometric-guided Self-Attention. Ultimately, the output of TopoFormer is a traffic topology scene graph, encapsulating the topological relationships among lanes, and guided by various road signals associated with the lanes.
@@
Traffic Topology Reasoning. STSU [4] introduced a lane topology reasoning approach to comprehend road structures, focusing on lane centerline detection from a Bird’s Eye View (BEV). The method comprises three stages: BEV feature construction, centerline detection, and centerline connection prediction. TPLR [5] and LaneGAP [28] subsequently refined the centerline representation, thereby enhancing its continuity and overall shape accuracy. The online map construction methodologies, such as MapTR [27], VectorMapNet [31], BeMapNet [40], and Gemap [61], proposed lane topology reasoning techniques grounded on geometric information, concurrently modeling diverse map elements like lane lines, pedestrian crossings, and curbs. To address more intricate road scenarios, TopoNet suggested the use of Graph Neural Networks (GNN) to forge meaningful connections between driving lanes and traffic signs. TopoMLP [53] and TopoLogits [13] amplified the inference capability of the topology structure by utilizing the lane’s spatial position. Our method also employs a geometric-guided Lane Aggregation Layer, but we introduce spatial information to more effectively illustrate the relationship between lanes, rather than augmenting the lane’s inherent features.
@@
Counterfactual Intervention. Counterfactual Intervention (CI) is widely applied in various reasoning tasks such as VQA [34, 3, 38], Re-ID [42, 25], and scene understanding [17, 45], etc. Lv et al. [34] model the physical knowledge relationships among different objects and apply them as counterfactual interventions to enhance the ability of causal reasoning. Similarly, Rao et al. [42] improve recognition capabilities in Re-ID tasks by introducing counterfactual interventions on the feature map and maximizing to encourage the network to learn more useful attention for the task. Inspired by these methods, we utilize the geometric information of lanes to construct the structure among lanes (i.e., road structure), and apply counterfactual interventions to this structure, maximizing the Total Indirect Effect (TIE) to encourage the model to learn more reasonable road structures for topology reasoning.
@@ 3.2 Lane Aggregation Layer
Geometry-guided Self-Attention (GSA) is proposed in the layer for the message passing in the graph, which is different from the conventional self-attention described in [46], as shown in Figure 2. Inspired by [13], the geometric distance among lane centerlines can serve as the basis for global dependencies, thereby enhancing the accuracy of lane topology reasoning. Therefore, we introduce the spatial proximity matrix (SPM) [62] to describe the normalized inverse geometric distances among lanes. It can be formulated as:
@@
For more detail about FFN please refer to [46]. It is noteworthy that different from TopoMLP [53] which introduces position encoding to enhance the features of each lane centerline, our proposed LAL explicitly utilizes position encoding for the aggregation of global information (i.e., spatial interaction among lanes), to exhibit stronger generalization.
@@ 3.5 Training and Inference
where $\lambda_{cls}$ and $\lambda_{reg}$ are the coefficients. The classification loss $\mathcal{L}_{cls}$ is a Focal loss [30] and the regression loss $\lambda_{reg}$ is an L1 loss. Note that the regression loss is calculated on the normalized 3D coordinates of the predicted ordered points list $\hat{\mathcal{V}}^{p}$. For the loss of edge prediction $\mathcal{L}_{\mathcal{E}}$, following [37], the total indirect effect of learned road structure on the prediction can be represented by the difference between the observed prediction ^$\hat{\mathcal{E}}_{A}=\hat{\mathcal{E}}(do(A=A),X=\widetilde{X})$ and its counterfactual result ^$\hat{\mathcal{E}}_{\overline{A}}=\hat{\mathcal{E}}(do(A=\overline{A}),X=\widetilde{X})$:
@@ 3.6 Traffic topology reasoning
where $Q_{out}^{T},Q_{init}^{T}\in\mathbb{R}^{N_{te}\times 256}$ represent the randomly initialized query and the output query from the last layer, respectively, where $N_{te}$ denotes the number of queries. The elements $\hat{T}^{b}=[\hat{t}_{i}^{b}\in\mathbb{R}^{4}\mid i=1,2,...,N_{te}]$ and $\hat{T}^{c}=[\hat{t}_{i}^{c}\in\mathcal{C}_{te}\mid i=1,2,...,N_{te}]$ correspond to the bounding box and class of the traffic elements, respectively. The TE Head is constructed by multilayer perceptron (MLPs). For the detection of the lane, we employ the detection method of $\text{T}^{2}\text{SG}$.
@@ 4.1 Datasets and Evaluation Setting
where $f$ is the square root function. More details can be found in the supplementary materials.
Figure 3: Qualitative results of the $\text{T}^{2}\text{SG}$ generation task and the lane topology reasoning, comparing the performance of TopoNet [23] and our proposed TopoFormer. The first row represents multi-view inputs. The second row illustrates the results of lane detection and lane topology reasoning. The third row visualizes our defined $\text{T}^{2}\text{SG}$, with TopoNet’s results converted to the same format for comparison. In these visualizations, the color green signifies correct predictions, red denotes erroneous predictions, and blue indicates missing predictions.
@@ 4.2 Implementation Details
We implement our model based on Pytorch [36] and MMdetection [9] on 16 Tesla V100 GPUs with a total batch size of 16. All images are resized into the same resolution of 1550 × 2048, and we use ResNet-50 [16] backbone pre-trained on ImageNet [12] paired with a Feature Pyramid Network (FPN) [29] to extract multi-scale features. The dimension of hidden feature $d$ is set to 256. The size of BEV grids is set to 200×200. For Lane centerline detection, the number of queries $N$ is 200, and the number of points in centerlines $l$ is 11. For the traffic elements detection, the number of queries is 100. The overall model TopoFormer is trained by AdamW optimizer [20] with a weight decay of 0.01. The learning rate is initialized with 2e-4 and we employ a cosine annealing schedule [32] for the learning rate.
@@ 4.3 Results
Traffic topology reasoning. We present the quantitative performance of our TopoFormer in traffic topology reasoning in Table 2, our method surpasses other methods with a 46.3% OLS in $subset_{A}$. Compared with TopoNet [23], a graph-based method, our method achieved higher score in the topology reasoning task (24.1 v.s. 10.9 on $\text{TOP}_{ll}$, 29.5 v.s. 23.8 on $\text{TOP}_{lt}$) while also achieves decent centerlines detection score (34.7 v.s. 28.5 on $\text{DET}_{l}$). This is due to the $\text{T}^{2}\text{SG}$ effectively capturing the global information among lanes by LAL, thereby enhancing the model’s performance in topology reasoning. Meanwhile, when compared to methods such as TopoLogic [13] that also utilize geometric information, our TopoFormer still demonstrates improvement on $\text{DET}_{l}$ and $\text{TOP}_{lt}$, owing to its effective modeling of reasonable road structure by CIL. Furthermore, our TopoFormer outperforms all other models across all metrics on $subset_{B}$, where the centerline is annotated in 2D space, showcasing its superior generalization capability

### v2

@@ Abstract
Understanding the traffic scenes and then generating high-definition (HD) maps present significant challenges in autonomous driving. In this paper, we defined a novel Traffic Topology Scene Graph ($\text{T}^{2}\text{SG}$), a unified scene graph explicitly modeling the lane, controlled and guided by different road signals (e.g., right turn), and topology relationships among them, which is always ignored by previous high-definition (HD) mapping methods. For the generation of $\text{T}^{2}\text{SG}$, we propose TopoFormer, a novel one-stage Topology Scene Graph TransFormer with two newly-designed layers. Specifically, TopoFormer incorporates a Lane Aggregation Layer (LAL) that leverages the geometric distance among the centerline of lanes to guide the aggregation of global information. Furthermore, we proposed a Counterfactual Intervention Layer (CIL) to model the reasonable road structure (e.g., intersection, straight) among lanes under counterfactual intervention. Then the generated $\text{T}^{2}\text{SG}$ can provide a more accurate and explainable description of the topological structure in traffic scenes. Experimental results demonstrate that TopoFormer outperforms existing methods on the $\text{T}^{2}\text{SG}$ generation task, and the generated $\text{T}^{2}\text{SG}$ significantly enhances traffic topology reasoning in downstream tasks, achieving a state-of-the-art performance of 46.3 OLS on the OpenLane-V2 benchmark. Our source code is available at https://github.com/MICLAB-BUPT/T2SG.
@@ 1 Introduction
Understanding the traffic scene is the key component of autonomous driving. Except for detecting and recognizing individual elements, the vehicles need to infer the topology relationship among them. Conventional traffic scene understanding tasks, such as lane perception [30], road signal elements detection [4], and high definition (HD) mapping [15] focus primarily on the isolated elements (i.e., map elements, and road signal elements ) but miss the relationship among them. To model the unified HD mapping for lane perception and road signal elements association, the traffic topology reasoning task on OpenLane-V2 [33] has recently been proposed. As a map-like reasoning results shown in Figure 1(c), the reasoning task aims to create a topology graph of the detected elements and thus facilitate decision-making in the downstream tasks, such as ego planning [5] and motion prediction [35, 45, 24].
@@
(2) We propose a Topology Scene Graph Transformer (TopoFormer) for $\text{T}^{2}\text{SG}$ task, which captures global dependencies among lanes with a Lane Aggeration Layer.
(3) We introduce a Counterfactual Intervention Layer to emphasize the reasonable road structure influencing lane connectivity and effectiveness in topology reasoning tasks.
(4) We evaluate our TopoFormer in scene graph generation task and show it outperforms all state-of-the-art methods. Furthermore, we attain a 46.3 OLS on the traffic topology reasoning benchmark, OpenLane-V2 [33], demonstrating the effectiveness of the proposed $\text{T}^{2}\text{SG}$ and TopoFormer for downstream tasks.
@@ 2 Related Work
Scene Graph Generation. Scene graphs were first introduced for Image Retrieval [11, 26], which broke down an image into its constituent objects, their attributes, and the relationships between them. The Visual Genome dataset [12] advanced image understanding, enabling scene graph extraction methods [6, 10, 25, 23]. Scene graphs are increasingly applied in tasks like image captioning [39] and visual question answering [34]. [32] introduced 3D scene graphs for object relationship perception, with subsequent research exploring GCN and Transformer-based methods [21] for point cloud registration [29] and scene generation [41]. Unlike prior approaches, $\text{T}^{2}\text{SG}$ pioneers scene graphs in traffic scene understanding, modeling lanes as nodes and their connections as edges.
@@
Figure 2: The overview of our proposed TopoFormer. Given the input multi-view images, we employ a DETR-like detector to identify lane objects with corresponding class and centerline coordinates. Subsequently, TopoFormer infers the relationships among these objects, which, along with the objects themselves, constitute the $\text{T}^{2}\text{SG}$. The main components of TopoFormer include two newly designed layers: (a) the Counterfactual Intervention Layer incorporating Counterfactual Self-Attention, and (b) the Lane Aggregation Layer incorporating Geometric-guided Self-Attention. Ultimately, the output of TopoFormer is a traffic topology scene graph, encapsulating the topological relationships among lanes, and guided by various road signals associated with the lanes.
@@
Traffic Topology Reasoning. STSU [2] introduced a lane topology reasoning approach for road structure comprehension, focusing on Bird’s Eye View (BEV) lane centerline detection through three stages: BEV feature construction, centerline detection, and connection prediction. TPLR [3] and LaneGAP [16] refined centerline representation, enhancing continuity and shape accuracy. Online map construction methods like MapTR [15], VectorMapNet [19], BeMapNet [27], and Gemap [43] integrated lane topology reasoning with geometric information, modeling map elements such as lane lines, pedestrian crossings, and curbs. TopoNet addressed complex road scenarios by using Graph Neural Networks to connect driving lanes and traffic signs. TopoSeq [40] proposed randomized prompt-to-sequence learning for joint extraction of lane topology from Directed Acyclic Graphs and geometric lane graphs. TopoMLP [37] and TopoLogits [7] enhanced topology inference by leveraging lane spatial positions. In contrast, our method employs a geometric-guided Lane Aggregation Layer, introducing spatial information to better illustrate lane relationships rather than augmenting lane features.
@@
Counterfactual Intervention. Counterfactual Intervention (CI) is widely used in reasoning tasks like VQA [22], Re-ID [28], and scene understanding [9].  [22] modeled the physical knowledge relationships among different objects and apply them as counterfactual interventions to boost causal reasoning, while  [28] enhanced Re-ID by applying CI to feature maps, maximizing task-relevant attention. Inspired by these, we leverage lane geometry to construct road structures and apply CI, maximizing the Total Indirect Effect (TIE) to encourage the model to learn more reasonable road structures for topology reasoning.
@@ 3.2 Lane Aggregation Layer
Geometry-guided Self-Attention (GSA) is proposed in the layer for the message passing in the graph, which is different from the conventional self-attention, as shown in Figure 2. Inspired by [7], the geometric distance among lane centerlines can serve as the basis for global dependencies, thereby enhancing the accuracy of lane topology reasoning. Therefore, we introduce the spatial proximity matrix (SPM) [44] to describe the normalized inverse geometric distances among lanes. It can be formulated as:
@@
It is noteworthy that different from TopoMLP [37] which introduces position encoding to enhance the features of each lane centerline, our proposed LAL explicitly utilizes position encoding for the aggregation of global information (i.e., spatial interaction among lanes), to exhibit stronger generalization.
@@ 3.5 Training and Inference
where $\lambda_{cls}$ and $\lambda_{reg}$ are the coefficients. The classification loss $\mathcal{L}_{cls}$ is a Focal loss [18] and the regression loss $\lambda_{reg}$ is an L1 loss. Note that the regression loss is calculated on the normalized 3D coordinates of the predicted ordered points list $\hat{\mathcal{V}}^{p}$. For the loss of edge prediction $\mathcal{L}_{\mathcal{E}}$, the total indirect effect of learned road structure on the prediction can be represented by the difference between the observed prediction ^$\hat{\mathcal{E}}_{A}=\hat{\mathcal{E}}(do(A=A),X=\widetilde{X})$ and its counterfactual result ^$\hat{\mathcal{E}}_{\overline{A}}=\hat{\mathcal{E}}(do(A=\overline{A}),X=\widetilde{X})$:
@@ 3.6 Traffic topology reasoning
where $Q_{out}^{T},Q_{init}^{T}\in\mathbb{R}^{N_{te}\times 256}$ represent the randomly initialized query and the output query from the last layer, respectively, where $N_{te}$ denotes the number of queries. The elements $\hat{T}^{b}=[\hat{t}_{i}^{b}\in\mathbb{R}^{4}\mid i=1,2,...,N_{te}]$ and $\hat{T}^{c}=[\hat{t}_{i}^{c}\in\mathcal{C}_{te}\mid i=1,2,...,N_{te}]$ correspond to the bounding box and class of the traffic elements, respectively. The TE Head is constructed by multilayer perceptron (MLPs). For lane detection, we employ the detection method of $\text{T}^{2}\text{SG}$.
@@ 4.1 Datasets and Evaluation Setting
where $f$ is the square root function.
Figure 3: Qualitative results of the $\text{T}^{2}\text{SG}$ generation task and the lane topology reasoning, comparing the performance of TopoNet [13] and our proposed TopoFormer. The first row represents multi-view inputs. The second row illustrates the results of lane detection and lane topology reasoning. The third row visualizes our defined $\text{T}^{2}\text{SG}$, with TopoNet’s results converted to the same format for comparison. In these visualizations, green signifies correct predictions, red denotes erroneous predictions, and blue indicates missing predictions.
@@ 4.2 Implementation Details
We implement our model based on Pytorch and MMdetection on 16 Tesla V100 GPUs with a total batch size of 16. All images are resized into the same resolution of 1550 × 2048, and we use ResNet-50 [8] backbone pre-trained on ImageNet paired with a Feature Pyramid Network (FPN) [17] to extract multi-scale features. The dimension of hidden feature $d$ is set to 256. The size of BEV grids is set to 200×100. For Lane centerline detection, the number of queries $N$ is 200, and the number of points in centerlines $l$ is 11. For the traffic elements detection, the number of queries is 100. The overall model TopoFormer is trained by AdamW optimizer with a weight decay of 0.01. The learning rate is initialized with 2e-4 and we employ a cosine annealing schedule [20] for the learning rate. Following [10], to accelerate convergence, we first train the lane detector and subsequently train the $\text{T}^{2}\text{SG}$ task using the pre-trained lane detector.
@@ 4.3 Results
Traffic topology reasoning. We present the quantitative performance of our TopoFormer in traffic topology reasoning in Table 2, our method surpasses other methods with a 46.3% OLS in $subset_{A}$. Compared with TopoNet [13], a graph-based method, our method achieved higher score in the topology reasoning task (24.1 v.s. 10.9 on $\text{TOP}_{ll}$, 29.5 v.s. 23.8 on $\text{TOP}_{lt}$) while also achieves decent centerlines detection score (34.7 v.s. 28.5 on $\text{DET}_{l}$). This is attributed to $\text{T}^{2}\text{SG}$’s ability to capture global lane information via LAL, enhancing topology reasoning performance. Compared to methods like TopoLogic [7], which also use geometric information, our TopoFormer improves $\text{DET}_{l}$ and $\text{TOP}_{lt}$ by effectively modeling road structure with CIL. Additionally, TopoFormer outperforms all models on $subset_{B}$, where centerlines are annotated in 2D space, demonstrating superior generalization.
@@ 5 Conclusion
6 Acknowledgement
This work is partly supported by the Funds for the National Natural Science Foundation of China under Grant 62202063 and U24B20176, Beijing Natural Science Foundation (L243027), Beijing Major Science and Technology Project under Contract No. Z231100007423014.